OpenStudy (anonymous):

Given the function f(x) = 3x, find the value of f−1(81). f−1(81) = 1 f−1(81) = 2 f−1(81) = 3 f−1(81) = 4

OpenStudy (anonymous):

@jhonyy9

OpenStudy (tennis5518):

you need to solve 3x=81

OpenStudy (triciaal):

do you mean 3^x?

OpenStudy (anonymous):

yep... my bad

OpenStudy (tennis5518):

if 3x=y and x=1/3y and finv tells you what x you put in given a resultant y so x=1/3y is your f^-1 y=81 x=1/3*81=27 i think

OpenStudy (soprano.h.d0816):

What do you think it is?

OpenStudy (triciaal):